A company produces steel rods. The lengths of the steel rods are normally distributed with a mean of 172.4-cm and a standard deviation of 1.4-cm. For shipment, 28 steel rods are bundled together.

Find the probability that the average length of a randomly selected bundle of steel rods is less than 172.9-cm.
P(M < 172.9-cm) =
Answer #1

Solution :

Given that ,

mean = = 172.4

standard deviation = = 1.4   

n = 28

= 172.4

=  / n = 1.4/ 28=0.26

P(M < 172.9) = P[(M - ) / < (172.9-172.4) /0.26 ]

= P(z < 1.92)

Using z table  

= 0.9726